Find meassures of the angles of each triangle
(2x+4)
(2x-9)
x

Answer by stanbon(48526)

Find measures of the angles of each triangle
(2x+4)
(2x-9)
x
---------
The sum of the angles of a triangle is 180 degrees.
-----------
EQuaTIOn:
x + 2x-8 + 2x+4 = 180
5x -4 = 180
5x = 184
x = 36.8 degrees (one of the angles)
2x-9 = 64.6 degrees ( 2nd angle)
2x+4 = 77.6 degrees ( 3rd angle)
===================================
Cheers,
Stan H.